OpenStudy (anonymous):

How would u do 17-48=3(x-3)-7

OpenStudy (shadowfiend):

So first off, what is 17 - 48?

OpenStudy (anonymous):

17-48= -31 3(x-3)-7 you do rainbow so 3 x (x) = 3x then 3(-3) = -9

OpenStudy (anonymous):

so -31=3x-9-7 so now you combine like terms u get -31=3x-16

OpenStudy (anonymous):

17-48 is -31

OpenStudy (anonymous):

now add 16 to both sides cancels the negative 16 so now its -15=3x

OpenStudy (anonymous):

now divide by 3 in both sides

OpenStudy (anonymous):

-15/3 = -5

OpenStudy (anonymous):

so x = -5

OpenStudy (shadowfiend):

needhelp4 nailed it :)

OpenStudy (anonymous):

shadow can u help me